OpenStudy (anonymous):

Write a direct variation equation that relates x and y. Assume that y varies directly as x. Then solve. If y = 1 when x = 4, find x when y = 8. a. y = x; c. y = x; b. y = x; d. y = x;

OpenStudy (anonymous):

a. y=1/4:33 b. y=-1/4x:-32 c. y=1/4x: 32 d. y=3/4x:32/3

OpenStudy (anonymous):

C

OpenStudy (anonymous):

hanks